OpenStudy (anonymous):

solve by eliminatin method 4x+5y=3, 3x-10y=6

OpenStudy (anonymous):

3x-10y=6 8x+10y=6 (multiplying eqn 1 by 2) --------- 11x=12 x=12/11 4(12/11)+5y=3 5y=3-(36/11)=-3/11 y=-3/55

OpenStudy (anonymous):

so there are infinitely many solutions correct

OpenStudy (anonymous):

no only one

OpenStudy (anonymous):

what the solution answer

OpenStudy (anonymous):

(12/11,-3/55)

OpenStudy (anonymous):

so the answer turns out to be a fraction

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

check the solution, if it satisfies the eqn

OpenStudy (anonymous):

i thought these were dependent equations - whenever you solve a system of equation and the anse. ends up being a numbe equals to same number a=a there are infinitely many solutions maybe i'm wrong

OpenStudy (anonymous):

oh the value of y is wrong, 5y=3-(48/11)=-15/11 y=-3/11

OpenStudy (anonymous):

not dependent

OpenStudy (anonymous):

(12/11,-3/11) is the solution